hi......can you please help me how to solve this problem.? "Steve is twice as old as Sylvie. Sylvie is three years older than Jacob. 4 years ago, Sylvie was twice Jacob's age. How old is Sylvie?"... I really tried to solve this for so many times but still i can't get the correct answer..please help me.
Answer by ankor@dixie-net.com(22740)

Let t = Steve's present age
Let s = Sylvie's present age
Let j = Jacob's present age
:
Write an equation for each statement:
:
Steve is twice as old as Sylvie.
t = 2s
:
Sylvie is three years older than Jacob.
s = j + 3
:
4 years ago, Sylvie was twice Jacob's age.
s - 4 = 2(j - 4)
s - 4 = 2j - 8
s = 2j - 8 + 4
s = 2j - 4
Replace s with (j+3), from the 2nd statement.
j + 3 = 2j - 4
3 + 4 = 2j - j
7 = j
:
Jacob is 7 yrs old
:
How old is Sylvie?
s = j + 3
s = 7 + 3
s = 10 yrs old is Sylvie
:
:
Confirm this in the statement:
"4 years ago, Sylvie was twice Jacob's age."
10 - 4 = 2(7 - 4)
6 = 2(3)
:
:
Poor Steve seems to have been totally left out, but he is 20 yrs old